OpenStudy (anonymous):

A parallelogram has an area of 224 square centimeters and a height of 14 centimeters. What is the length of the base of the parallelogram? cm

OpenStudy (anonymous):

ya can u help me

jimthompson5910 (jim_thompson5910):

Area = base * height A = b*h 224 = b*14 b = ???

OpenStudy (anonymous):

16

OpenStudy (anonymous):

224/14=16 224=16*14 b=16

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

can uhelp me again

OpenStudy (anonymous):

yes or no

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

ya